Required information [The following information applies to the questions displayed below.] Marin Company produces two software products (Cloud-X and Cloud-Y) in two separate departments (A and B). These products are highly regarded network maintenance programs. Cloud-X is used for small networks and Cloud-Y is used for large networks. Marin is known for the quality of its products and its ability to meet dates promised for software upgrades. Department A produces Cloud-X, and department B produces Cloud-Y. The production departments are supported by two support departments, systems design and programming services. The sources and uses of the support department time are summarized as follows: From To Total Labor Hours Design Programming Department A Department B Design - 5,000 1,000 9,000 15,000 Programming 400 - 600 1,000 2,000 The costs in the two service departments are as follows: Design Programming Labor and materials (all variable) $ 50,000 $ 36,000 Depreciation and other fixed costs 40,000 4,000 Total $ 90,000 $ 40,000 Required: 1. Determine the total support costs allocated to each of the producing departments using (a) the direct method, (b) the step method (design department goes first), and (c) the reciprocal method? (Do not round intermediate calculations. Round your final answer to nearest whole dollar amount.)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Marin
Direct Method Design Programming Dept. A Dept. B Total
Design Hours               9,000.00         15,000.00         24,000.00
Design Hours % 37.50% 62.50% 100.00%
Programming Hours               1,000.00            2,000.00            3,000.00
Programming Hours % 33.33% 66.67% 100.00%
Total costs          90,000.00      40,000.00       130,000.00
Design cost to be allocated- Design Hour Feet %        (90,000.00)                     -               33,750.00         56,250.00                        -  
Programming cost to be allocated- Programming Hour %     (40,000.00)             13,333.00         26,667.00                        -  
Total cost after allocation                        -                       -               47,083.00         82,917.00       130,000.00
Step Method Design Programming Dept. A Dept. B Total
Design Hours        1,000.00               9,000.00         15,000.00         25,000.00
Design Hours % 4.00% 36.00% 60.00% 100.0000%
Programming Hours               1,000.00            2,000.00            3,000.00
Programming Hours % 33.33% 66.67% 100.00%
Total costs          90,000.00      40,000.00       130,000.00
Design cost to be allocated- Design Hour Feet %        (90,000.00)        3,600.00             32,400.00         54,000.00                        -  
Total costs after allocation of design costs                        -        43,600.00             32,400.00         54,000.00       130,000.00
Programming cost to be allocated- Programming Hour %     (43,600.00)             14,533.00         29,067.00                        -  
Total cost after allocation                        -                       -               46,933.00         83,067.00       130,000.00
Reciprocal Method Design Programming Dept. A Dept. B Total
Design Hours                        -          1,000.00               9,000.00         15,000.00         25,000.00
Programming Hours               400.00                     -                 1,000.00            2,000.00            3,400.00
Design Hours % 0.00% 4.00% 36.00% 60.00% 100.00%
Programming Hours % 11.76% 0.00% 29.41% 58.82% 100.00%
Total costs          90,000.00      40,000.00       130,000.00
Let cost of Design is $ X and Programming is $ Y
So, X= 90000+ 0.1176Y
Y= 40000+ 0.04X
Substituting the value of X in equation 2:
Y = 40000 + 0.04(90000 + 0.1176Y)
Y = 40000 + 3600+.0047Y
Y – 0.0047Y = 40000+3600
0.9953Y = 43,600
Y = 43600/0.9953
Y = $ 43,806
Substituting the value of Y in equation 1:
So, X= 90000+ .1176Y
So, X= 90000+ 0.1176*43806
So, X= 90000+ 5154
X= $ 95,154
Allocation Summary Design Programming Dept. A Dept. B Total
Design Hours %                        -   4.00% 36.00% 60.00% 100.00%
Programming Hours % 11.76%                     -   29.41% 58.82% 100.00%
Department cost          90,000.00      40,000.00                            -                          -         130,000.00
Total costs
Design        (95,154.00)        3,806.00             34,255.00         57,093.00                        -  
Programming            5,154.00     (43,806.00)             12,884.00         25,768.00                        -  
Cost after allocation                        -                       -               47,139.00         82,861.00       130,000.00